OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6 i choose c but i got it wrong

OpenStudy (anonymous):

the given point does not belong to the line you've chosen

OpenStudy (anonymous):

can u show me

OpenStudy (anonymous):

the line is of the form y=ax+b since this line is parallel to y=x+8 then it has the same slope so a=1=>y=x+b now (4;10) belongs to this line =>10=4+b=>b=6 so the answer is y=x+6 which isB.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw